OpenStudy (anonymous):

The total amount of weight on an elevator cannot exceed 1,800 pounds. A delivery person is on the elevator delivering packages. Together, the delivery person and the packages weigh 520 pounds. If the average weight of a person is 160 pounds, set up and solve an inequality to determine how many additional people, p, the elevator can hold.

OpenStudy (fanduekisses):

\[520+160p \le1800\]

OpenStudy (fanduekisses):

I think ;)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@pgpilot326

OpenStudy (fanduekisses):

Yep I just checked, and I am pretty sure this is right ;)

OpenStudy (fanduekisses):

if you solve the inequality, you'd find out that the elevator can hold 8 more people

OpenStudy (anonymous):

but these are the answer choices @Fanduekisses

OpenStudy (fanduekisses):

yeah, it's a. it's the same, p=8 (p= the amount of people the elevator will hold)

OpenStudy (anonymous):

so would it be a b c or d lol i feel stupid but im just having a rough day sorry please? @Fanduekisses

OpenStudy (fanduekisses):

It's alright, I understand lol I've felt the exact same way ;p it's A.

OpenStudy (anonymous):

\[p \le \frac{ 1800-520 }{160 }\]

OpenStudy (anonymous):

thanks!!!

OpenStudy (fanduekisses):

just like 520+160p≤1800 (except that in A. the 520 and 160p are in like different positions, but it's the same, 520+160p<=1800 and 160p+520<=1800

OpenStudy (fanduekisses):

No problem ;)

OpenStudy (anonymous):

oooh, the commutative property in action!